# First Resolution of Microlensed Images of a Binary-Lens Event

### v1

@@ 2.3 HERMES Spectroscopic Data & Stellar Parameters
We obtained a high-resolution optical spectrum (sobject_id 220125001601051) on UT 2022-01-25.75 as part of ongoing observations of the Galactic Archaeology with HERMES survey (De Silva et al., 2015, hereafter GALAH). The total exposure time is 3600 s. The HERMES spectrograph on the 3.9-m Anglo-Australian Telescope at Siding Spring Observatory covers four wavelength regions (4713–4903, 5648–5873, 6478–6737, and 7585–7887 Å) that can show absorption features from up to 31 elements, including the two strong Balmer lines $\mathrm{H\alpha}$ and $\mathrm{H\beta}$.
@@
We use the observations and analysis of the most recent analysis cycle, which will be published as the fourth GALAH data release (DR4; Buder et al., in preparation). The spectra are reduced using a similar analysis to the third data release (Buder et al., 2021) with a custom-built reduction pipeline (Kos et al., 2017). Stellar parameters, that is, effective temperature ($T_{\mathrm{eff}}$), surface gravity ($\log g$), radial, microturbulence and rotational velocities ($v_{\mathrm{rad}}$, $v_{\mathrm{mic}}$ and $v\sin i$), as well as up to 31 elemental abundances, are then simultaneously estimated by minimizing the weighted residuals between the observed and synthetic stellar spectra. The latter are interpolated with a high-dimensional neural network (compare to Ting et al., 2019) trained on synthetic stellar spectra generated with the spectrum synthesis code Spectroscopy Made Easy (Piskunov & Valenti, 2017), and degraded to the measured instrumental resolution of the spectrum. This change with respect to the third data release, which used wavelength regions of carefully selected unblended lines, allows 94% of the observed wavelength range to be used and thus considerably increases the measurement precision.
@@
The spectroscopic fit also includes photometric and astrometric information from the 2MASS (Skrutskie et al., 2006) and Gaia surveys (Lindegren et al., 2021) to constrain the surface gravity. This leads to more accurate stellar parameters, as shown in the previous HERMES analyses (Buder et al., 2018, 2019, 2021), particularly for cool giant stars such as the target of this study, where molecular features complicate the spectroscopic analysis. The fit is very good with the signal-to-noise ratio increasing across the four wavelength regions (16, 107, 194, and 254), as can be seen from Figure 2. We do not expect a good fit to the Balmer line cores or the lithium lines at $4861$, $6563$, and $6708\,\text{\AA}$, respectively. No quality concerns are raised by the automatic pipeline (flag_sp = 0).
@@
The final radial velocity of $v_{\mathrm{rad}}=-72.44\pm 0.18\,\mathrm{km\,s^{-1}}$ is in very good agreement with the value from the Gaia DR3 spectrum, that is, $v_{\mathrm{rad},Gaia}=-71.8\pm 1.0\,\mathrm{km\,s^{-1}}$ (Recio-Blanco et al., 2023). The final stellar parameters are $T_{\mathrm{eff}}=3776\pm 67\,\mathrm{K}$, $\log g=1.05\pm 0.12$, $\mathrm{[Fe/H]}=-0.148\pm 0.053$, $v_{\mathrm{mic}}=1.82\pm 0.28\,\mathrm{km\,s^{-1}}$, and $v\sin i=5.5\pm 1.4\,\mathrm{km\,s^{-1}}$. The stellar abundances are close to the Solar values, for example, $\mathrm{[Mg/Fe]}=0.007\pm 0.011$.
@@ 3.1 Single-lens Model
where $A(t)$ is the magnification as a function of time, and $f_{\rm S}$ and and $f_{\rm B}$ are the source flux and the blended flux within the PSF, respectively. In the simplest single-lens single-source (1L1S) model, i.e., the PSPL model (Paczyński, 1986), the magnification is given by,
@@
The peak of the light curve (Figure 1) is flattened compared to the PSPL model, suggesting possible finite-source effects (Gould, 1994a). We fit the finite-source point-lens (FSPL) model to the light curve by introducing an additional parameter $\rho={\theta_{*}}/{\theta_{\rm E}}$. In the FSPL model, $A_{\rm FSPL}$ is calculated by integrating the $A_{\rm PSPL}$ over a limb-darkened source disk. Using the best-fit spectroscopic parameters (see § 2.3), we estimate the limb-darkening coefficients $(\Gamma_{g^{\prime}},\Gamma_{r^{\prime}},\Gamma_{i^{\prime}},\Gamma_{G})=(0.91,0.77,0.63,0.74)$ based on Claret & Bloemen (2011) and Claret (2019).
The best-fit FSPL model is shown as a dashed black line in the top panel of Figure 1, and the residuals are shown in the bottom panel. The residuals exhibit significant trends deviating from the FSPL model, spanning $\sim 10$ days around the peak. These systematic deviations cannot be absorbed by adjusting either the limb-darkening coefficients or the free parameters in the FSPL models.
@@ 3.2 Binary-lens Model
The light curve can be well described by the binary-lens single-source (2L1S) model. We start the analysis with the static binary-lens model, which includes three additional parameters $(s,q,\alpha)$ to describe a non-rotating binary-lens systems: $s$ is the projected angular separation of the binary components normalized by $\theta_{\rm E}$, $q$ is the binary mass ratio, and $\alpha$ is the angle between the source-lens trajectory and the binary-lens axis. We use the advanced contour integration package VBBinaryLensing (Bozza, 2010; Bozza et al., 2018) to calculate the binary-lens magnification $A(t,\Theta_{\rm 2L1S})$ given the parameters $\Theta_{\rm 2L1S}=\{s,q,\alpha,\rho,t_{0},u_{0},t_{\rm E}\}$.
@@
We start by scanning the 7-d parameter space on a fixed $(\log s,\log q,\alpha)$ grid with $-1\leq\log s\leq 1$, $-2\leq\log q\leq 1$ and $0\leq\alpha<2\pi$. At each grid point, we minimize the $\chi^{2}$ by setting the 4 remaining parameters $(t_{0},u_{0},t_{\rm eff}\equiv u_{0}t_{\rm E},t_{*}\equiv\rho t_{\rm E})$ free in MCMC using the EMCEE package (Foreman-Mackey et al., 2013). The resulting $\chi^{2}$ map is shown in Figure 3, from which we identify 5 local minima in the close-binary $(s<1)$ regime and 9 local minima in the wide-binary $(s>1)$ regime. The severe finite-source effects smear out the sharp magnification structure of caustics, leading to a plethora of local minima over a broad range of $\log q$. For close solutions, there is a perfect degeneracy of $(\log q,\alpha)\to(-\log q,\alpha+\pi)$, and therefore we only conduct the grid search with $\log q\leq 0$. For wide solutions, we have a 4-fold degeneracy in $\alpha$. When $q\gg 1$, the “planetary” caustic associated with secondary mass is reduced to a Chang-Refsdal (Chang & Refsdal, 1979) caustic with shear $s^{-2}/q\ll 1$. The four identical cusps of a Chang-Refsdal lead to a degeneracy of $\alpha\to\alpha+\pi/2$, as shown in Figure 3. The solutions with $q\gg 1$ can correspond to unrealistic value of $t_{\rm E}$. We limited our analysis to $\log q<3$, which corresponds to an upper limit of $t_{\rm E}\sim 2000\,\rm d$.
@@
Then we probe all local minima by setting all the parameters free in MCMC. We also introduce the higher-order effects into the modeling. We include the annual microlens parallax effects (Gould, 1992), which are the light-curve distortions due to the Earth’s orbital acceleration toward the Sun. The static binary-lens model has a perfect geometric degeneracy $(u_{0},\alpha)\to-(u_{0},\alpha)$, which is broken when taking the microlens parallax into account. For the $u_{0}>0$ and $u_{0}<0$ geometry, the lens takes different sides of the source trajectory, leading to different magnifications. The microlens parallax effects are described by ${\bm{\pi}}_{\rm E}$ vector with the East and North components ${\bm{\pi}}_{\rm E}=(\pi_{\rm E,E},\pi_{\rm E,N})$, as defined in Gould (2004). We also consider the orbital motion of the binary lens, which can have degeneracies with microlens parallax (Batista et al., 2011; Skowron et al., 2011). We apply the linearized orbital motion
@@
with two parameters $ds/dt$ and $d\alpha/dt$. This approximation is subject to physical constraints as discussed in Dong et al. (2009). For a bound system, the ratio of the projected kinetic to the potential energy $\beta$ should be less than unity:
@@
We find that the best-fit solution is a wide-binary with $u_{0}<0$ and label it as “wide A--”. The source trajectory and caustics are shown in the upper left sub-panel of Figure 4. The “wide A--” solution has $\chi^{2}=948.9$, and all other solutions are worse by $\Delta\chi^{2}>50$. Nevertheless, we keep all solutions with $\Delta\chi^{2}<81$ (listed in Table 1) for further analysis incorporating the VLTI GRAVITY data.
@@ 4.1 Parameterization
The images generated from the binary-lens equation are oriented with respect to the binary-lens axis and in the units of the Einstein radius, so they need to be rotated to the sky plane and scaled by $\theta_{\rm E}$ to compute the closure phases. We define $\Psi$ as the position angle of the binary-lens axis. The direction of lens-source relative motion ${\bm{\mu}}_{\rm rel}$ , which is the same as that of the microlens parallax ${\bm{\pi}}_{\rm E}$, can be expressed as:
@@
We incorporate the 2 parameters $\{\theta_{\rm E},\Psi\}$ in addition to the 2L1S parameters $\{s,q,\alpha,\rho,t_{0},u_{0},t_{\rm E}\}$ for modeling the interferometric data. The microlens parallax is re-parameterized by $(\pi_{\rm E,N},\pi_{\rm E,E})=\pi_{\rm E}\cdot(\cos\Phi_{{\bm{\pi}}},\sin\Phi_{{\bm{\pi}}})$ with $\Phi_{{\bm{\pi}}}=\alpha+\Psi$.
@@ 4.2 Probing Interferometric Observables
Prior to carrying out the full joint analysis of VLTI and light curve, we first investigate the VLTI observables with limited input from the light-curve models. We start the investigation using only $(s,q,\rho)$ from the best-fit light-curve model “wide A--”. We perform 4-dimensional grid search on $(\theta_{\rm E},\Psi,u_{x},u_{y})$ to fit the closure-phase data from each of the eight VLTI exposures, where $(u_{x},u_{y})$ are the source’s coordinates on the binary-lens plane. At each grid point, we generate the images scaled and rotated to the sky plane based on $(\theta_{\rm E},\Psi)$ and then calculate the closure phases and corresponding $\chi_{{\rm VLTI},i}^{2}(\theta_{\rm E},\Psi,u_{x},u_{y})$ for every VLTI exposure $i$. We describe our method of calculating the closure phase from the binary-lens images in Appendix B. Our closure-phase calculations from the images are validated with the PMOIRED software [^5] https://github.com/amerand/PMOIRED (Mérand, 2022).
To exclude unreasonable source positions, we add a penalty term
@@
for the deviation from light-curve magnification for each VLTI exposure $i$, where $A_{{\rm LC},i}$ is the magnification from the “wide A--” model at the $i$-th VLTI exposure, and we adopt a fractional error of $10\%$ so that $(u_{x,i},u_{y,i})$ are only subject to loose light-curve magnification constraints.
@@
The left panel of Figure 5 shows the $\Delta\chi^{2}_{\rm tot}(\theta_{\rm E},\Psi)$ map, from which we identify four local minima. The $\theta_{\rm E}$ is within a broad range of $0.6-1.0\,{\rm mas}$, and $\Psi$ is subject to a 4-fold degeneracy around $(-100^{\circ},-10^{\circ},80^{\circ},170^{\circ})$, each with an uncertainty of $\sim 10^{\circ}$. The 4-fold degeneracy corresponds to having the source positions near the four cusps of the “central” caustic associated with the primary mass, resulting in similar image morphologies that are compatible with the VLTI data (see the sub-panels on the right of Figure 5).
Next, we further probe the four local minima by jointly fitting all epochs of VLTI data. In a satisfactory VLTI microlensing model, the source positions should lie approximately on a straight line following the source-lens relative motion, so we parameterize the source positions $(u_{x},u_{y})$ with $(u_{0},t_{0},t_{\rm E},\alpha)$ as,
@@
where $\tau=(t-t_{0})/{t_{\rm E}}$. The interferometric parameters $(\theta_{\rm E},\Psi)$ are seeded from the above-mentioned four local minima. We allow $(\log s,\log q,\log\rho)$ to freely vary, and no longer add the $\chi^{2}$ penalty term as in Eq. 9. The four best-fit solutions have similar $\chi^{2}$ ($\Delta\chi^{2}<10$), so the VLTI data alone cannot distinguish between them. All the solutions strongly prefer $u_{0}<0$. As in the previous single-lens events, $u_{0}>0$ and $u_{0}<0$ solutions can distinguished with two or more epochs of VLTI data (Dong et al., 2019; Zang et al., 2020; Cassan et al., 2022). The four degenerate solutions correspond to 4 distinct $\alpha$ values, denoted as $\alpha_{\rm VLTI}$, which are displayed as shaded regions in Figure 6. The allowed $(\log s,\log q)$ from the posteriors have broad ranges, suggesting that the resulting constraints on $\alpha_{\rm VLTI}$ are not specific to the seed solution of “wide A--”, so we use them as the starting point for the joint analysis with the light curves in § 4.3.

### v3

@@ 2.3 HERMES Spectroscopic Data and Stellar Parameters
We obtained a high-resolution optical spectrum (sobject_id 220125001601051) on UT 2022-01-25.75 as part of ongoing observations of the Galactic Archaeology with HERMES survey (De Silva et al., 2015). The total exposure time is 3600 s. The HERMES spectrograph on the 3.9 m AAT at Siding Spring Observatory covers four wavelength regions (4713–4903, 5648–5873, 6478–6737, and 7585–7887 Å) that can show absorption features from up to 31 elements, including the two strong Balmer lines $\mathrm{H\alpha}$ and $\mathrm{H\beta}$.
@@
We use the observations and analysis of the most recent analysis cycle, which will be published as the fourth GALAH data release (Buder et al. 2024). The spectra are reduced using a similar analysis as that for the third data release (Buder et al., 2021), with a custom-built reduction pipeline (Kos et al., 2017). Stellar parameters—that is, effective temperature ($T_{\mathrm{eff}}$), surface gravity ($\log g$), and radial, microturbulence, and rotational velocities ($v_{\mathrm{rad}}$, $v_{\mathrm{mic}}$ and $v\sin i$), as well as up to 31 elemental abundances—are then simultaneously estimated by minimizing the weighted residuals between the observed and synthetic stellar spectra. The latter are interpolated with a high-dimensional neural network (compare to Ting et al., 2019), trained on synthetic stellar spectra generated with the spectrum synthesis code Spectroscopy Made Easy (Piskunov & Valenti, 2017), and degraded to the measured instrumental resolution of the spectrum. This change with respect to the third data release, which used wavelength regions of carefully selected unblended lines, allows 94% of the observed wavelength range to be used and thus considerably increases the measurement precision.
@@
The spectroscopic fit also includes photometric and astrometric information from the 2MASS (Skrutskie et al., 2006) and Gaia surveys (Lindegren et al., 2021a) to constrain the surface gravity. This leads to more accurate stellar parameters, as shown in the previous HERMES analyses (Buder et al., 2018, 2019, 2021), particularly for cool giant stars such as the target of this study, where molecular features complicate the spectroscopic analysis. The fit is very good, with the signal-to-noise ratio increasing across the four wavelength regions (16, 107, 194, and 254), as can be seen from Figure 2. We do not expect a good fit to the Balmer-line cores or the lithium lines at $4861$, $6563$, and $6708\,\text{\AA}$, respectively. No quality concerns are raised by the automatic pipeline (flag_sp = 0).
@@
The final radial velocity of $v_{\mathrm{rad}}=-72.44\pm 0.18\,\mathrm{km\,s^{-1}}$ is in very good agreement with the value from the Gaia DR3 spectrum—that is, $v_{\mathrm{rad,Gaia}}=-71.8\pm 1.0\,\mathrm{km\,s^{-1}}$ (Recio-Blanco et al., 2023). The final stellar parameters are $T_{\mathrm{eff}}=3776\pm 67\,\mathrm{K}$, $\log g=1.05\pm 0.12$, $\mathrm{[Fe/H]}=-0.148\pm 0.053$, $v_{\mathrm{mic}}=1.82\pm 0.28\,\mathrm{km\,s^{-1}}$, and $v\sin i=5.5\pm 1.4\,\mathrm{km\,s^{-1}}$. The stellar abundances are close to the solar values; for example, $\mathrm{[Mg/Fe]}=0.007\pm 0.011$.
@@ 3.1 Single-lens Model
where $A(t)$ is the magnification as a function of time, and $f_{\rm S}$ and $f_{\rm B}$ are the source flux and the blended flux within the PSF, respectively. In the simplest single-lens single-source (1L1S) model, i.e., the PSPL model (Paczyński, 1986), the magnification is given by
@@
The peak of the light curve (Figure 1) is flattened compared to the PSPL model, suggesting possible finite-source effects (Gould, 1994a). We fit the FSPL model to the light curve by introducing an additional parameter, $\rho={\theta_{*}}/{\theta_{\rm E}}$. In the FSPL model, $A_{\rm FSPL}$ is calculated by integrating $A_{\rm PSPL}$ over a limb-darkened source disk. Using the best-fit spectroscopic parameters (see § 2.3), we estimate the limb-darkening coefficients $(\Gamma_{g^{\prime}},\Gamma_{r^{\prime}},\Gamma_{i^{\prime}},\Gamma_{G},\Gamma_{K})=(0.91,0.77,0.63,0.74,0.29)$ based on Claret & Bloemen (2011) and Claret (2019).
The best-fit FSPL model is shown as a dashed black line in the top panel of Figure 1, and the residuals are shown in the bottom panel. The residuals exhibit significant trends deviating from the FSPL model, spanning $\sim 10$ days around the peak (at $\rm HJD^{\prime}\sim 9600$). These systematic deviations cannot be absorbed by adjusting either the limb-darkening coefficients or the free parameters in the FSPL models. Deviations from single-lens models suggest the possible existence of a companion to the lens or the source. There is also a $\sim 0.1$ mag broad bump-like feature in the Gaia $G$-band light curve at around $\sim 300$ days before the peak (see the right panel of Figure 3). Such an additional feature excludes single-lens binary-source models, and we study binary-lens models in the following.
@@ 3.2 Binary-lens Model
The light curve can be well described by the binary-lens single-source (2L1S) model. We start the analysis with the static binary-lens model, which includes three additional parameters $(s,q,\alpha)$ to describe a nonrotating binary-lens system: $s$ is the projected angular separation of the binary components normalized by $\theta_{\rm E}$, $q$ is the binary mass ratio, and $\alpha$ is the angle between the source-lens trajectory and the binary-lens axis. We use the advanced contour integration package VBBinaryLensing (Bozza, 2010; Bozza et al., 2018) to calculate the binary-lens magnification $A(t,\Theta_{\rm 2L1S})$, given the parameters $\Theta_{\rm 2L1S}=\{s,q,\alpha,\rho,t_{0},u_{0},t_{\rm E}\}$.
@@
We start by scanning the seven-dimensional parameter space on a fixed $(\log s,\log q,\alpha)$ grid with $-1\leq\log s\leq 1$, $-2\leq\log q\leq 1$ and $0\leq\alpha<2\pi$. At each grid point, we minimize the $\chi^{2}$ by setting the four remaining parameters $(t_{0},u_{0},t_{\rm eff}\equiv u_{0}t_{\rm E},t_{*}\equiv\rho t_{\rm E})$ free in Markov Chain Monte Carlo (MCMC) using the EMCEE package (Foreman-Mackey et al., 2013). The resulting $\chi^{2}$ map is shown in Figure 4, from which we identify five local minima in the close-binary $(s<1)$ regime and nine local minima in the wide-binary $(s>1)$ regime. The severe finite-source effects smear out the sharp magnification structure of caustics, leading to a plethora of local minima over a broad range of $\log q$. For close solutions, there is a perfect degeneracy of $(\log q,\alpha)\to(-\log q,\alpha+\pi)$, and therefore we only conduct the grid search with $\log q\leq 0$. For wide solutions, we have a fourfold degeneracy in $\alpha$. When $q\gg 1$, the “planetary” caustic associated with secondary mass is reduced to a Chang-Refsdal (Chang & Refsdal, 1979) caustic with shear $s^{-2}/q\ll 1$. The four identical cusps of Chang-Refsdal lead to a degeneracy of $\alpha\to\alpha+\pi/2$, as shown in Figure 4. The solutions with $q\gg 1$ can correspond to an unrealistic value of $t_{\rm E}$. We limit our analysis to $\log q<3$, which corresponds to an upper limit of $t_{\rm E}\sim 2000\,$days.
@@
Then we probe all local minima by setting all the parameters free in MCMC. We also introduce higher-order effects into the modeling. We include the annual microlens parallax effects (Gould, 1992), which are the light-curve distortions due to the Earth’s orbital acceleration toward the Sun. The static binary-lens model has a perfect geometric degeneracy $(u_{0},\alpha)\to-(u_{0},\alpha)$, which is broken when taking the microlens parallax into account. For the $u_{0}>0$ and $u_{0}<0$ geometry, the lens takes different sides of the source trajectory, leading to different magnifications. As defined in Gould (2004), the microlens parallax vector ${\bm{\pi}}_{\rm E}$ is described by the east and north components ${\bm{\pi}}_{\rm E}=(\pi_{\rm E,E},\pi_{\rm E,N})$. We also consider the orbital motion of the binary lens, which can have degeneracies with microlens parallax (Batista et al., 2011; Skowron et al., 2011). We apply the linearized orbital motion
@@
with the two parameters $ds/dt$ and $d\alpha/dt$. This approximation is subject to physical constraints, as discussed in Dong et al. (2009). For a bound system, the ratio of the projected kinetic to the potential energy $\beta$ should be less than unity:
@@
We find that the best-fit solution is a wide binary with $u_{0}<0$ and we label it as “wide A--”. The source trajectory and caustics are shown in the upper left subpanel of Figure 3. The “wide A--” solution has $\chi^{2}=948.9$, and all other solutions are worse by $\Delta\chi^{2}>50$. Nevertheless, we keep all solutions with $\Delta\chi^{2}<81$ (listed in Table 1) for further analysis incorporating the VLTI GRAVITY data.
@@ 4.1 Parameterization
The images generated from the binary-lens equation are oriented with respect to the binary-lens axis and in the units of the Einstein radius, so they need to be rotated to the sky plane and scaled by $\theta_{\rm E}$ to compute the closure phases. We define $\Psi$ as the position angle of the binary-lens axis. The direction of the lens-source relative motion ${\bm{\mu}}_{\rm rel}$ , which is the same as that of the microlens parallax ${\bm{\pi}}_{\rm E}$, can be expressed as
@@
We incorporate the two parameters $\{\theta_{\rm E},\Psi\}$ in addition to the 2L1S parameters $\{s,q,\alpha,\rho,t_{0},u_{0},t_{\rm E}\}$ for modeling the interferometric data. The microlens parallax is reparameterized by $(\pi_{\rm E,N},\pi_{\rm E,E})=\pi_{\rm E}\cdot(\cos\Phi_{{\bm{\pi}}},\sin\Phi_{{\bm{\pi}}})$ with $\Phi_{{\bm{\pi}}}=\alpha+\Psi$.
@@ 4.2 Probing Interferometric Observables
Prior to carrying out the full joint analysis of VLTI and light curve, we first investigate the VLTI parameters with limited input from the light-curve models. We start the investigation using only $(s,q,\rho)$ from the best-fit light-curve model “wide A--.” We perform a four-dimensional grid search on $(\theta_{\rm E},\Psi,u_{x},u_{y})$ to fit the closure-phase data from each of the eight VLTI exposures, where $(u_{x},u_{y})$ are the source’s coordinates on the binary-lens plane. At each grid point, we generate the images scaled and rotated to the sky plane based on $(\theta_{\rm E},\Psi)$ and then calculate the closure phases and corresponding $\chi_{{\rm VLTI},i}^{2}(\theta_{\rm E},\Psi,u_{x},u_{y})$ for every VLTI exposure $i$. We describe our method of calculating the closure phase from the binary-lens images in Appendix B. Our closure-phase calculations from the images are validated with the PMOIRED software [^5] https://github.com/amerand/PMOIRED (Mérand, 2022).
To exclude unreasonable source positions, we add a penalty term,
@@
for the deviation from light-curve magnification for each VLTI exposure $i$, where $A_{{\rm LC},i}$ is the magnification from the “wide A--” model at the $i$th VLTI exposure, and we adopt a fractional error of $10\%$ so that $(u_{x,i},u_{y,i})$ are only subject to loose light-curve magnification constraints.
@@
The left panel of Figure 5 shows the $\Delta\chi^{2}_{\rm tot}(\theta_{\rm E},\Psi)$ map, from which we identify four local minima. The $\theta_{\rm E}$ is within a broad range of $0.6-1.0\,{\rm mas}$, and $\Psi$ is subject to a fourfold degeneracy around $(-100^{\circ},-10^{\circ},80^{\circ},170^{\circ})$, each with an uncertainty of $\sim 10^{\circ}$. The fourfold degeneracy corresponds to having the source positions near the four cusps of the “central” caustic associated with the primary mass, resulting in similar image morphologies that are compatible with the VLTI data (see the subpanels on the right of Figure 5).
Next, we further probe the four local minima by jointly fitting all epochs of VLTI data. In a satisfactory VLTI microlensing model, the source positions should lie approximately on a straight line, following the source-lens relative motion, so we parameterize the source positions $(u_{x},u_{y})$ with $(u_{0},t_{0},t_{\rm E},\alpha)$ as
@@
where $\tau=(t-t_{0})/{t_{\rm E}}$. The interferometric parameters $(\theta_{\rm E},\Psi)$ are seeded from the abovementioned four local minima. We allow $(\log s,\log q,\log\rho)$ to freely vary, and we no longer add the $\chi^{2}$ penalty term as in Eq. 9. The four best-fit solutions have similar $\chi^{2}$ ($\Delta\chi^{2}<10$), so the VLTI data alone cannot distinguish between them. All the solutions strongly prefer $u_{0}<0$. As in the previous single-lens events, $u_{0}>0$ and $u_{0}<0$ solutions can be distinguished with two or more epochs of VLTI data (Dong et al., 2019; Zang et al., 2020; Cassan et al., 2022). The four degenerate solutions correspond to four distinct $\alpha$ values, denoted as $\alpha_{\rm VLTI}$, which are displayed as the shaded regions in Figure 6. The allowed $(\log s,\log q)$ from the posteriors have broad ranges, suggesting that the resulting constraints on $\alpha_{\rm VLTI}$ are not specific to the seed solution of “wide A--,” so we use them as the starting point for the joint analysis with the light curves in § 4.3.